There is a 2% chance that a customer’s Visa charge will be rejected at a certain Target store because the transaction exceeds the customer’s credit limit. What is the probability that the first such rejection occurs on the third Visa transaction? Select one: a. 1.88% b. 98.12% c. 1.92% d. 98.08%
Answer #1

Option C is correct

Given that 0.02 is a chance that a customer’s Visa charge will be rejected at a certain Target store.

Thus the probability of non-rejection of card=1-0.02=0.98

Probability( Rejection occurs on third Visa transaction)= P(no reject in first)*P(no reject in second)* P(reject in 3)

= 0.98*0.98*0.02

=0.019208

In percentage 0.019208*100=1.9208

Answer: 1.92%